# Fast Feature Matching of UAV Images via Matrix Band Reduction-based GPU Data Schedule

### v1

@@ Abstract
Feature matching dominats the time costs in structure from motion (SfM). The primary contribution of this study is a GPU data schedule algorithm for efficient feature matching of Unmanned aerial vehicle (UAV) images. The core idea is to divide the whole dataset into blocks based on the matrix band reduction (MBR) and achieve efficient feature matching via GPU-accelerated cascade hashing. First, match pairs are selected by using an image retrieval technique, which converts images into global descriptors and searches high-dimension nearest neighbors with graph indexing. Second, compact image blocks are iteratively generated from a MBR-based data schedule strategy, which exploits image connections to avoid redundant data IO (input/output) burden and increases the usage of GPU computing power. Third, guided by the generated image blocks, feature matching is executed sequentially within the framework of GPU-accelerated cascade hashing, and initial candidate matches are refined by combining a local geometric constraint and RANSAC-based global verification. For further performance improvement, these two seps are designed to execute parallelly in GPU and CPU. Finally, the performance of the proposed solution is evaluated by using large-scale UAV datasets. The results demonstrate that it increases the efficiency of feature matching with speedup ratios ranging from 77.0 to 100.0 compared with KD-Tree based matching methods, and achieves comparable accuracy in relative and absolute bundle adjustment (BA). The proposed algorithm is an efficient solution for feature matching of UAV images.
@@
[label3]organization=Engineering Research Center of Natural Resource Information Management and Digital Twin Engineering Software, Ministry of Education,city=Wuhan, postcode=430074, country=China
@@
[label4]organization=State Key Laboratory of Information Engineering in Surveying, Mapping and Remote Sensing, Wuhan University,city=Wuhan, postcode=430072, country=China
@@ 1 Introduction
Unmanned aerial vehicle (UAV) has been increasingly becoming a critical remote sensing platform in photogrammetry and computer vision (Jiang et al., 2021b; Yao et al., 2019), and it has been widely used in a range of applications, e.g., indoor inspection of transmission lines (Jiang et al., 2017), procedural 3D modeling of urban facilities (Wang et al., 2023), and precision management of agricultural plants (Zheng et al., 2021). The precise calculation of camera poses is an essential preliminary step, which has been usually implemented through structure from motion (SfM) (Jiang et al., 2020a). In the workflow of SfM, the time costs are dominated by feature matching due to the quadratic combination complexity of match pairs and nearest neighbor searching (NNS) of high-dimension feature descriptors. Thus, fast feature matching is required for recent SfM systems.
@@
In the literature, extensive work has been reported to address these issues. To decrease the combination complexity of match pairs, auxiliary data and visual similarity are the widely used clues to select overlapped image pairs and avoid exhaustive feature matching (Hartmann et al., 2016), e.g., the onboard POS (Positioning and Orientation System) data from UAV platforms (Jiang and Jiang, 2020; Remondino and Gerke, 2015) and the geometric priors from data acquisitions (Jiang et al., 2024; Schönberger et al., 2014). In addition, to cope with oblique data acquisition and geometry-aware trajectory planning (Li et al., 2023), visual similarity has also been extensively exploited in modern SfM systems, which usually use pre-trained visual words to encode images into fix-length vectors and convert the problem of selecting overlapped image pairs into the one of nearest neighbor searching among the vectors, such as BoW (Bag-of-Words) (Nister and Stewenius, 2006; Havlena and Schindler, 2014; Jiang et al., 2022a) and VLAD (Vector of Locally Aggregated Descriptors) (Hou et al., 2023; Jiang et al., 2023). By using these techniques, the time complexity of feature matching is significantly decreased, which is transited from a quadratic relationship to a linear one with respect to the number of involved UAV images.
@@
Match pair selection reduces the combination complexity of images. However, high time costs are still required in the exhaustive nearest neighbor searching among two sets of high-dimension feature descriptors (Hartmann et al., 2016), especially for the widely used KD-Tree based SIFT feature matching (Lowe, 2004). To accelerate feature matching, existing methods can be divided into two groups, i.e., redesign of feature descriptors and acceleration of feature matching. For the former, feature descriptors are designed to decrease the computational burden, such as the ORB (Rublee et al., 2011) with binary descriptors. However, as verified in recent work, SIFT-based feature matching is still the most robust algorithm in practical engineering applications (Ji et al., 2023; Jiang et al., 2021a). For the latter, the hardware acceleration technique has been exploited to speed up feature matching. The classical examples include the SIFTGPU (Wu, 2011) and popsift (Griwodz et al., 2018) algorithms that exploit GPU (Graphic Processing Unit) for feature matching. In contrast to the KD-Tree based approximate nearest neighbor (ANN) searching, hashing-based methods have also gained lots of attention in feature matching, whose core idea is to convert the real-value descriptors to discrete binary codes via hashing functions and map similar descriptors of query and database images into the same cell (Cheng et al., 2014). To achieve higher speedup ratios, GPU acceleration and data schedule have been integrated into hashing-based matching pipelines in the recent work (Xu et al., 2017; Zhang et al., 2023), which use hardware acceleration for calculating hash codes and similarity scores and exploit data schedule to relief input/output (IO) burdens.
@@
Compared with classical SIFT matching algorithms, hashing-based methods can achieve promising speedup ratios for feature matching of UAV images. However, some issues still exist in the proposed solutions. First, the performance of existing data schedule depends on the connection structures between images. In the work of Xu et al. (2017), the group-block based data schedule strategy assumes dense image connections. For oblique photogrammetry, image connections are usually very sparse, and it degenerates the overall performance of GPU-based hashing matching. Second, existing data schedule strategies emphasize more on the IO burdens, however, ignoring the parallel computing power of modern GPU. For the data schedule used in Zhang et al. (2023), image loading and free sequences are first generated based on the image connection structure, and feature matching is executed sequentially according to loading and free sequence. In other words, only one image is matched against other images in GPU memory. Third, outliers are prone to retain in feature matching due to the usage of binary hashing codes for calculating descriptor similarity scores. It would cause high time cost in classical RANSAC (Random Sample Consensus) based outlier removal (Lu et al., 2016). Thus, this study proposes a matrix band reduction-based GPU data schedule for fast feature matching of UAV images. Our main contributions are summarized as follows: (1) we propose a matrix band reduction (MBR) based data schedule strategy to divide the sparsely connected view graph into compact blocks, which can adapt well to the connection structure of images and the memory volume of GPUs; (2) we design a cascade hashing based feature matching workflow that integrates the spatial angular order (SAO) based local constraint and RANSAC-based global verification for outlier removal. Combined with MBR-based data schedule, the proposed image matching solution can increase the usage of GPU for feature matching and CPU for outlier removal; (3) we verify the performance of the proposed feature matching solution by using large-scale datasets and compare it with state-of-the-art software packages.
@@
This paper is organized as follows. Section 2 presents the workflow of the GPU data schedule algorithm for efficient feature matching of UAV images. 3 conducts the performance evaluation and comparison with other methods. Finally, Section 4 presents the conclusions.
@@ 2 Methodology
This study proposes a fast feature matching of UAV images via matrix band reduction-based GPU data schedule as shown in Figure 1. There are three major steps, i.e., view graph construction, schedule block generation, and cascade hashing feature matching. First, SIFT features are extracted from each image and used to calculate VLAD-based image descriptors using a pre-trained codebook. After indexed in the HNSW (Hierarchical Navigable Small World) based graph structure, overlapped image pairs are retrieved and used to construct the view graph. Second, the initial view graph is first compressed through a matrix band reduction algorithm, which permutes the image order and ensures that overlapped match pairs are located near the diagonal region in the view graph. Data schedule blocks are then created from the permuted adjacent matrix. Through the iterative execution of MBR, a list of schedule blocks can be generated. Finally, images in each schedule block are loaded into GPU memory for calculating hashing codes and generating initial matching results. At the same time, outlier removal is simultaneously conducted on the CPU by using spatial angular order-based local geometric constraint and RANSAC-based global geometric constraint. The details are presented in the following sections.
Figure 1: The workflow of the proposed algorithm.
2.1 Overview of cascade hashing feature matching
Figure 2: The principle of cascade hashing feature matching.
(2) Hashing remapping for fine encoding. $m$ bits coarse hashing increases the recall of feature matching while the precision is low. Thus, candidate matches are remapped into $n\ (n>m)$ bits long binary codes with LSH. The Hamming distance that ranges from 00 to $n$ is computed between the query point $p$ and their candidate matches, and it is used to sort candidate matches ascendly.
2.2 Matrix band reduction-based GPU data schedule for feature matching
GPU-accelerated cascade hashing gains extremely high efficiency. However, the hard disk to GPU data schedule and limited volume of GPU memory degenerates its overall performance. Although promising efficiency has been reported in recent studies (Xu et al., 2017; Zhang et al., 2023), these methods either rely on dense image connections or not fully exploit the computing power of modern GPU. Therefore, this study proposes a matrix band reduction-based GPU data schedule for fast feature matching, which aims to adapt to both densely and sparsely connected images and use GPU computing power as high as possible.
2.2.1 View graph construction based on image retrieval
@@
To address this challenge, the study employs a high-efficiency image retrieval framework (Jiang et al., 2023). The proposed methodology integrates two core components: (1) the VLAD representation for compact image encoding, and (2) the HNSW graph structure for efficient vector indexing. The VLAD descriptor utilizes a minimal-size codebook to aggregate local feature descriptors into a unified high-dimensional global representation, significantly reducing computational redundancy. Concurrently, the HNSW index architecture accelerates the approximate nearest neighbor search process over the encoded VLAD vectors, enabling rapid similarity matching while maintaining sublinear time complexity. This synergistic combination balances representational compactness with scalable search performance.
@@
2.2.2 Schedule block generation via iterative matrix band reduction
Data schedule between hard disk and GPU is required since all feature descriptors cannot be loaded into GPU with limited memory at one time. In general, existing strategies for data schedule can be grouped into three categories. The most direct strategy is to execute feature matching sequentially for all selected match pairs. This strategy, however, causes frequent data load and free in GPU. For the second category, data load and free lists are first generated according to their connections, and feature matching is conducted by the interleaved execution of reading in data in the load list and clearing out data in the free list. This strategy decreases data schedule burdens. However, it does not fully exploit the computing power of GPU since only the newly loaded image is matched with other images in GPU. For the third category, images are divided into blocks that can be loaded into GPU at one time, and feature matching is then executed between loaded images. It has high efficiency in pair-wise feature matching. For sparsely connected images, its performance decreases obviously.
@@ 2.2.2 Schedule block generation via iterative matrix band reduction
(2) Calculate permute order. By using the MBR algorithm, a permute order ${order}_{perm}$ is generated, which permutes the row and column index of $M_{sym}$ and creates the MBR matrix $M_{mbr}$. This study adopts the Gibbs-Poole-Stockmeyer (GPS) algorithm (Gibbs et al., 1976) due to its low time costs and memory requirements. In addition, the image id list ${list}_{id}$ is updated according to ${order}_{perm}$.
(3) Create new schedule blocks. By using MBR matrix $M_{mbr}$, new schedule blocks are created under the limitation of the GPU memory. As presented in Figure 3(b), the GPU memory size ${Size}_{gpu}$ is restricted by the yellow grids, which indicates the maximum number of descriptor data that can be loaded into GPU memory. By defining the block size ${Size}_{blk}$, new schedule blocks ${{block}_{ij}}$ are created as rendered by red rectangles in Figure 3(b), in which $i$ and $j$ represents the row and column index of each block, respectively. Block $\{{block}_{ij}\}$ stores its corresponding match pairs.
@@
(5) Termination. Iteratively executing steps (2)-(4) until the MBR matrix $M_{mbr}$ is empty. Finally, the schedule blocks $\{{block}_{ij}\}$ can be obtained.
Compared with existing data schedule strategies, the proposed solution has two major advantages. First, it can minimize data load and free in GPU as much as possible because of the usage of MBR algorithm for adjacent matric compression and block generation; second, a high usage ratio of GPU can be achieved because match pairs as many as possible are packed into one schedule block.
2.2.3 Cascade hashing matching enhanced by geometric constraints
Cascade hashing feature matching is then implemented using the generated data schedule blocks $\{{block}_{ij}\}$. As the sum of data volume in each block row $i$ is restricted by GPU memory, feature descriptors in the block row can be loaded into GPU at one time. Supposing that generated blocks $\{{block}_{ij}\}$ consists of $k$ rows and $l$ columns, i.e., $i=1,2,\ldots,k$ and $j=1,2,\ldots,l$, and the current block row is represented as $\{{block}_{i1},{block}_{i2},\ldots,{block}_{il}\}$. The pipeline of cascade feature matching is implemented as follows:

### v2

# Efficient Feature Matching for UAV Images based on Compact GPU Data Scheduling

@@ Abstract
Feature matching dominates the time costs in structure from motion (SfM). The primary contribution of this study is a GPU data schedule algorithm for efficient feature matching of Unmanned aerial vehicle (UAV) images. The core idea is to divide the whole dataset into blocks based on matrix band reduction (MBR) and achieve efficient feature matching via GPU-accelerated cascade hashing. First, match pairs are selected by using an image retrieval technique, which converts images into global descriptors and searches high-dimension nearest neighbors with graph indexing. Second, compact image blocks are iteratively generated from a MBR-based data schedule strategy, which exploits image connections to generate image blocks and increase the usage of GPU computing power. Third, guided by the generated image blocks, feature matching is executed sequentially within the framework of GPU-accelerated cascade hashing, and initial candidate matches are refined by combining a local geometric constraint and RANSAC-based global verification. For further performance improvement, these two steps are designed to execute in parallel in GPU and CPU. Finally, the performance of the proposed solution is evaluated by using large-scale UAV datasets. The results demonstrate that it increases the efficiency of feature matching with speedup ratios ranging from 77.0 to 100.0 compared with KD-Tree based matching methods due to its high usage of GPU computing power. Besides, it achieves comparable accuracy in both relative and absolute bundle adjustment (BA). The proposed algorithm is an efficient solution for feature matching of large-scale UAV images.
@@
[label3]organization=Shenzhen Key Laboratory of Spatial Smart Sensing and Services,city=Shenzhen, postcode=518060, country=China
@@
[label4]organization=School of Surveying and Mapping, Information Engineering University,city=Zhengzhou, postcode=450001, country=China
[…]
[label5]organization=Guangdong Laboratory of Artificial Intelligence and Digital Economy (Shenzhen),city=Shenzhen, postcode=518060, country=China
@@ 1 Introduction
Unmanned aerial vehicle (UAV) has been increasingly becoming a critical remote sensing platform in photogrammetry and computer vision (Jiang et al., 2021b; Yao et al., 2019), and it has been widely used in a range of applications, e.g., indoor inspection of transmission lines (Jiang et al., 2017), procedural 3D modeling of urban facilities (Wang et al., 2023), and precision management of agricultural plants (Zheng et al., 2021). The precise calculation of camera poses is an essential preliminary step, which has been usually implemented through structure from motion (SfM) (Jiang et al., 2020a). In the workflow of SfM, the time costs are dominated by feature matching due to the quadratic combination complexity of match pairs and nearest neighbor search (NNS) of high-dimension feature descriptors. Therefore, fast feature matching is required for recent SfM systems.
Figure 1: GPU data schedule algorithm for feature matching and SfM reconstruction.
@@
In the literature, extensive work has been reported to address these issues. To decrease the combination complexity of match pairs, auxiliary data and visual similarity are the widely used clues to select overlapped image pairs and avoid exhaustive feature matching (Hartmann et al., 2016), e.g., the onboard POS (Positioning and Orientation System) data from UAV platforms (Jiang and Jiang, 2020; Remondino and Gerke, 2015) and the geometric priors from data acquisitions (Jiang et al., 2024b; Schönberger et al., 2014). In addition, to cope with oblique data acquisition and geometry-aware trajectory planning (Li et al., 2023), visual similarity has also been extensively exploited in modern SfM systems, which usually use pre-trained visual words to encode images into fix-length vectors and convert the problem of selecting overlapped image pairs into the one of nearest neighbor search among the vectors, such as BoW (Bag-of-Words) (Nister and Stewenius, 2006; Havlena and Schindler, 2014; Jiang et al., 2022a) and VLAD (Vector of Locally Aggregated Descriptors) (Hou et al., 2023; Jiang et al., 2023). By using these techniques, the time complexity of feature matching is significantly decreased, which is transited from a quadratic relationship to a linear one with respect to the number of involved UAV images.
@@
Match pair selection reduces the combination complexity of images. However, high time costs are still required in the exhaustive nearest neighbor search among two sets of high-dimension feature descriptors (Hartmann et al., 2016), especially for the widely used KD-Tree based SIFT feature matching (Lowe, 2004). Nowadays, deep learning based methods have also been reported to achieve reliable feature matching, such as SuperGlue (Sarlin et al., 2020) and LoFTR (Sun et al., 2021). Although they avoid nearest neighbor searching, the localization precision still the main drawback for their usage in SfM. To accelerate feature matching, existing methods can be divided into two groups, i.e., redesign of feature descriptors and acceleration of feature matching. For the former, feature descriptors are designed to decrease the computational burden, such as the ORB (Rublee et al., 2011) with binary descriptors. However, as verified in recent work, SIFT-based feature matching is still the most robust algorithm in practical engineering applications (Ji et al., 2023; Jiang et al., 2021a). For the latter, the hardware acceleration technique has been exploited to speed up feature matching. The classical examples include the SIFTGPU (Wu, 2011) and popsift (Griwodz et al., 2018) algorithms that exploit GPU (Graphic Processing Unit) for feature matching. In contrast to the KD-Tree based nearest neighbor search, hashing-based methods have also gained lots of attention in feature matching, whose core idea is to convert the real-value descriptors to discrete binary codes via hashing functions and map similar descriptors of query and database images into the same cell (Cheng et al., 2014). To achieve higher speedup ratios, GPU acceleration and data schedule have been integrated into hashing-based matching pipelines in the recent work (Xu et al., 2017; Zhang et al., 2023), which use hardware acceleration for calculating hash codes and similarity scores and exploit data schedule to relief input/output (IO) burdens.
@@
Compared with classical SIFT matching algorithms, hashing-based methods can achieve promising speedup ratios for feature matching of UAV images. However, some issues still exist in the proposed solutions. First, the performance of existing data schedule depends on dense image connections. Pair-wise feature matching is first executed for all images to create an image adjacency matrix, and the group-block based data schedule strategy is used to divide the whole dataset into groups and blocks that are fed into GPU sequentially for feature matching (Xu et al., 2017). However, for oblique photogrammetry, image connections are usually very sparse, and it degenerates the overall performance of GPU-based hashing matching since a majority of blocks have few image pairs. Second, existing data schedule strategies emphasize more on the IO burdens, however, ignoring the parallel computing power of modern GPU. For the data schedule used, image loading and free sequences are first generated based on the image connection structure, and feature matching is executed sequentially according to loading and free sequence (Zhang et al., 2023). In other words, only one image is matched against other images in GPU memory. Third, outliers are prone to retain in feature matching due to the usage of binary hashing codes for calculating descriptor similarity scores. It would cause high time cost in classical RANSAC (Random Sample Consensus) based outlier removal (Lu et al., 2016).
To address above-mentioned issues, this study proposes a matrix band reduction-based GPU data schedule for fast feature matching of UAV images, as shown in Figure 1. Our main contributions are summarized as follows: (1) we propose a matrix band reduction (MBR) based data schedule strategy to divide the sparsely connected view graph into compact blocks, which can adapt well to the connection structure of images and the memory volume of GPUs; (2) we design a cascade hashing based feature matching workflow that integrates the spatial angular order (SAO) based local constraint and RANSAC-based global verification for outlier removal. Combined with MBR-based data schedule, the proposed image matching solution can increase the usage of GPU for feature matching and CPU for outlier removal; (3) we verify the performance of the proposed feature matching solution using large-scale datasets and compare it with state-of-the-art software packages.
@@
This paper is organized as follows. Section 2 presents the workflow of the GPU data schedule algorithm for efficient feature matching of UAV images. Section 3 conducts the performance evaluation and comparison with other methods. Finally, Section 5 presents the conclusions.
Figure 2: The workflow of the proposed algorithm.
@@ 2 Methodology
This study proposes a fast feature matching solution for UAV images via matrix band reduction-based GPU data schedule, as shown in Figure 2. There are three major steps, i.e., view graph construction, schedule block generation, and cascade hashing feature matching. First, SIFT features are extracted from each image and used to calculate VLAD-based image descriptors using a pre-trained codebook. After indexed in the HNSW (Hierarchical Navigable Small World) based graph structure, overlapped image pairs are retrieved and used to construct the view graph. Second, the initial view graph is first compressed through the matrix band reduction algorithm, which permutes the image order and ensures that overlapped match pairs are located near the diagonal region in the view graph. Data schedule blocks are then created from the permuted adjacent matrix. Through the iterative execution of MBR, a list of schedule blocks can be generated. Finally, images in each schedule block are loaded into GPU memory for calculating hashing codes and generating initial matching results. At the same time, outlier removal is simultaneously conducted on the CPU by using spatial angular order-based local geometric constraint and RANSAC-based global geometric constraint. The details are presented in the following sections.
2.1 View graph construction based on image retrieval
@@
Figure 3: The workflow of the VLAD-HNSW based image retrieval workflow.
To address this challenge, the study employs a high-efficiency image retrieval framework (Jiang et al., 2023), as shown in Figure 3. The proposed method includes three steps: (1) codebook generation based on image and feature selection; (2) the VLAD representation for compact image encoding; and (3) the HNSW graph structure for efficient vector indexing. VLAD descriptors utilize a minimal-size codebook to aggregate local feature descriptors into a unified high-dimensional global representation, significantly reducing computational redundancy. The HNSW index architecture accelerates the approximate nearest neighbor search process over the encoded VLAD vectors, enabling rapid similarity matching while maintaining sublinear time complexity. This synergistic combination balances representational compactness with scalable search performance.
@@
2.2 Schedule block generation via matrix band reduction
Data schedule between hard disk and GPU is required because all feature descriptors cannot be loaded into GPU with limited memory at one time. In general, existing strategies for data schedule can be grouped into three categories, as illustrated in Figure 4. The most direct strategy is to execute feature matching sequentially for all selected match pairs. This strategy, however, causes frequent data load and free in GPU. For the second category, data load and free lists are first generated according to their connections, and feature matching is conducted by the interleaved execution of reading in data in the load list and clearing out data in the free list. This strategy decreases data schedule burdens. However, it does not fully exploit the computing power of GPU since only the newly loaded image is matched with other images in GPU. For the third category, images are divided into blocks that can be loaded into GPU at one time, and feature matching is then executed between loaded images. It has high efficiency in pair-wise feature matching. For sparsely connected images, its performance decreases obviously.
Figure 4: The illustration of existing data schedule strategies between hard disk and GPU memory.
@@ 2.2 Schedule block generation via matrix band reduction
(2) Calculate permute order. By using the MBR algorithm, a permute order ${order}_{perm}$ is generated, which permutes the row and column index of $M_{sym}$ and creates the MBR matrix $M_{mbr}$. This study adopts the Gibbs-Poole-Stockmeyer (GPS) algorithm (Gibbs et al., 1976) due to its low time costs and memory requirements. For more details, refering to Appendex A. In addition, the image id list ${list}_{id}$ is updated according to ${order}_{perm}$.
(3) Create new schedule blocks. By using MBR matrix $M_{mbr}$, new schedule blocks are created under the limitation of the GPU memory. As presented in Figure 5(b), the GPU memory size ${Size}_{gpu}$ is restricted by the yellow grids, which indicates the maximum number of descriptor data that can be loaded into GPU memory. By defining the block size ${Size}_{blk}$, new schedule blocks ${block}_{ij}$} are created as rendered by red rectangles in Figure 5(b), in which $i$ and $j$ represents the row and column index of each block, respectively. Block $\{{block}_{ij}\}$ stores its corresponding match pairs. For each block ${block}_{ij}$, images with their indices within the corresponding red rectangle are included, and their descirptor data would be loaded into GPU memory for feature matching.
@@
(5) Termination. Iteratively executing steps (2)-(4) until the MBR matrix $M_{mbr}$ is empty.
Compared with existing strategies, the proposed data schedule solution has two advantages. First, it minimizes data load and free in GPU as much as possible due to the usage of MBR for adjacent matric compression and block generation. Second, a high usage ratio of GPU can be achieved since match pairs as many as possible are packed into one schedule block.
2.3 Cascade hashing-based feature matching
2.3.1 The workflow of cascade hashing feature matching
Figure 6: The workflow of cascade hashing-based feature matching.
(2) Hashing remapping for fine encoding. $m$ bits coarse hashing increases the recall of feature matching while the precision is low. Thus, candidate matches are remapped into n(n>m)n\ (n>m) bits long binary codes with LSH. The Hamming distance that ranges from 0 to $n$ is computed between the query point $p$ and their candidate matches, and it is used to sort candidate matches ascendly.
2.3.2 Schedule block guided cascade hashing feature matching
Based on the principle of cascade hashing feature matching, the proposed solution is then implemented using the generated data schedule blocks $\{{block}_{ij}\}$. As the sum of data volume in each block row $i$ is restricted by GPU memory, feature descriptors in the block row can be loaded into GPU at one time. Supposing that generated blocks $\{{block}_{ij}\}$ consists of $k$ rows and $l$ columns, i.e., $i=1,2,\ldots,k$ and $j=1,2,\ldots,l$, and the current block row $i$ is represented as $\{{block}_{i1},{block}_{i2},\ldots,{block}_{il}\}$. The pipeline of cascade feature matching is implemented as follows: